In a small country near Baltic sea there are three commodities


1
Expert's answer
2022-01-10T09:54:09-0500

Explanation:


Let's start by writing down all the components of the problem:

1. Potato's sacs ( P ) cost 2 crowns, denote the price of a potato sack by Pp 

2. Meatballs ( M ) cost 4 per crock, denote Pm as the price of meatballs

3. Jam cost 6 per jar ( J ), denoted Pj as the price of jam.

4. Gunnar has an Income: M=360


His budget constrain is then:


  • The amount he spends in potatoes "P_p \times P"
  • The amount he spends in meatballs "P_m \times M"
  • The amount he spends in jam "P_j \times J"


He only spends money on those goods, then his expenditures equal his income:


"I=P_p \times P+P_m \times M +P_j \times J"

"360=2 \times P +4 \times M +6 \times J"


(b) Next we need to re-express all prices so relative prices are the same as before:


If the new price of potatoes is "P^{'}_p=1," then the price of meatballs will be: "P^{'}_m=\frac{P_m}{P_p}=\frac{4}{2}=2"


(c) the same can be done for jam:


If the new price of potatoes is "P^{'}_p=1," then the price of meatballs will be: "P^{'}_m=\frac{P_m}{P_p}=\frac{6}{2}=3"


(d) Gunnar's Income would be than half as before: "I^{'}=\frac{I}{P_p}=\frac{360}{2}=180"



(e) We can summarize everything re-expressing Gunnars budget constraint

The old budget constraint was: "I = P_p \times P +P_m \times M + P_j \times J"


Now setting "P^{'}_p=1," is the same as dividing everything by "P_p=2"


"\frac{I}{Pp}=\frac{P_p}{P_p}\times P+\frac{P_m}{P_p}\times M + \frac{P_j}{P_p}\times J"


"I^{'}=P+P^{'}_m\times M +P^{'}_j\times J"


"180 = 1 \times P +2 \times M + 3 \times J"


Answer:


(a) "360 = 2 \times P +4 \times M +6 \times J"

(b) 2

(c) 3

(d) 180

(e) "180 = 1 \times P +2 \times M + 3 \times J"